OpenStudy (anonymous):

What is the slope of a line passing through (3, 4) and (5, 8)?

OpenStudy (hba):

m=y2-y1/x2-x1 = 5-3/8-4=2/4=1/2

OpenStudy (anonymous):

thanx

OpenStudy (hba):

ur welcome